OpenStudy (anonymous):

A rectangular pond is 21 ft squared in area. If the owner can surround the pond with a 20-foot fence, what are the dimensions of the pond?

OpenStudy (anonymous):

You have two equations. Area = l*w = 21 Perimeter = l + l + w + w = 20 re-arrange the area equation to solve for l or w, (l=21/w) 21/w + 21/w + w + w = 20 solve for l and w